a photon of yellow light of frequency
5.0 * 10^14Hz is emitted in a time of
10^-8 seconds
evaluate 1. the length of the wave train
2. the number of complete oscillations present in the wave train
1
Expert's answer
2021-10-20T10:39:02-0400

1) We can find the length of the wave train as follows:


"l=c\Delta t=3\times10^8\ \dfrac{m}{s}\times10^{-8}\ s=3.0\ m."

2) Let's first find the length of the wave:


"\lambda=\dfrac{c}{f}=\dfrac{3\times10^8\ \dfrac{m}{s}}{5.0\times10^{14}\ Hz}=6.0\times10^{-7}\ m."


We can find the number of complete oscillations present in the wave train as follows:


"N=\dfrac{l}{\lambda}=\dfrac{3.0\ m}{6.0\times10^{-7}\ m}=5.0\times10^6."